# A First Look at GPT Apps: Landscape and Vulnerability The original version was uploaded on Feb 23, 2024

### v2

@@ 5. GPT Apps Plagiarism in the Wild
The GPT app store serves as the main interface for users to retrieve GPT apps. As depicted in Figure 15, when users search for a specific keyword, OpenAI’s GPT store displays up to ten related search results including GPT app names, descriptions, conversation counts, and more. This section explores the impact of vulnerabilities in GPT apps on both the GPT app store and user experience, with a particular focus on the prevalence of plagiarism among these GPT apps in the wild. Specifically, we aim to answer two questions. First, are there similar GPT apps in terms of their names and descriptions that may affect user experience during keyword searches? Second, what are the similarities in system prompts, and why might GPT app creators be motivated to maintain such similarities?
@@ 5.1. Similarities in Names and Descriptions
We analyze duplicate and similar GPT apps in the GPT app store by searching for specific keywords. For this purpose, we created a search bot that automatically interacts with the GPT app in OpenAI GPT app store. It uses GPT app names from the top 10,000 GPT apps and records the results. Then, we extracted all relevant GPT apps from the results and verified: (1) whether the targeted GPT app and each GPT app in search results share identical names and descriptions; and (2) whether the target GPT app and the matched GPT app have the same name and semantically similar descriptions. To measure semantic similarity, we transformed the descriptions of two GPT apps to be evaluated into embeddings using a BERT model (bert-base-multilingual-uncased) (Wolf et al., 2020) and computed their cosine similarity. If the value is larger than 0.99, we treat them are similar.
@@
We first analyzed the number of GPT apps sharing both the same name and description as the targeted GPT app. The findings, depicted in Figure 16 (left), reveal that 135 GPT apps have an identical counterpart; 21 out of 10,000 GPT apps have two identical matches. A minor portion of GPT apps have complete duplication beyond two. However, the criterion requiring GPT apps to have exactly identical names and descriptions is stringent. By loosening the criteria to include semantically similar descriptions, a larger number of similar GPT apps are identified. Figure 16 (right) shows that over 700 GPT apps have at least one similar counterpart. These outcomes suggest that users searching for a specific GPT app by name might encounter multiple GPT apps with identical names and closely related descriptions. This outcome may confuse users’ choices, as they can only see limited information on the search interface of OpenAI’s GPT app store.
@@
We further examined the conversation counts of these similar GPT apps, which is the primary indicator of popularity in the official GPT app store. Users often base their selection on the perceived popularity of GPT apps when faced with multiple similar options. As shown in Figure 17, we assessed the absolute difference in conversation counts between the targeted GPT app and its counterparts, finding that a smaller difference often complicates the decision-making process for users. Concretely, the cumulative distribution of these differences reveals that over 75% of GPT apps with semantically similar counterparts have a nearest equivalent with less than approximately 100 conversation count differences. This could significantly challenge users in choosing the appropriate GPT app. In particular, they may encounter GPT apps that severely affect their experience, such as those containing hazardous content or those that do not function as expected. Furthermore, the reliability of conversation counts as a metric is questionable (for example, it can be manipulated through repeated, non-substantive interactions), and the issue of GPT app plagiarism may aggravate the problem for users.
@@ 5.2. Similarities in System Prompts
Top 10,000 GPT apps. In this section, we first analyze the similarities between GPT apps beyond their names and descriptions. We aim to identify GPT apps with nearly identical system prompts among the top 10,000 GPT apps. We employ the same method for assessing prompt similarity as described in §$\S$5.1, using cosine similarity thresholds of 0.9 and 0.95.
@@
Figure 18. Similarity Analysis in System Prompts among the Top 10,000 GPT Apps. Cosine similarity bound: 0.9 (left) and 0.95 (right)
Figure 18 displays the number of GPT apps with 1–6 similar counterparts. In summary, over 91 GPT apps have a counterpart with system prompts showing a cosine similarity threshold of 0.9. With a stricter threshold of 0.95, more than 45 GPT apps still have potential duplicated counterparts, including 30 GPT apps with one potential duplicate each. We also validate the metadata of these GPT apps and find that none share the same names. In contrast to our observations in §$\S$5.1 that there exists a large amount of GPT apps having similar names and descriptions, few creators replicate GPT apps to resemble existing ones.
We generally categorize the similarities in system prompts of these GPT apps into two types. The first type consists of GPT apps created by the same creator with only minor modifications in their names and system prompts, tailored to serve slightly different application scenarios. For example, two GPT apps by the same creator differ by only one word; one translates input to English, while the other translates to Chinese. This likely stems from the lack of user-friendly functionality switching options on the interaction UI, as currently, user interactions with GPT apps are text-based. The majority of cases fall into the second type, where GPT apps are created by different authors and have distinct names. Although these creators vary, their objectives might be similar: either to create honeypots to attract all related traffic for this type of GPT app or to use leaked or open-sourced system prompts to build identical GPT apps for traffic diversion.
@@
Figure 19. Similarity Analysis in System Prompts among the Top 500 GPT Apps. “Success” targets GPT apps successfully in GPT configuration extraction, whereas “Fail” targets on rest of the GPT apps.
@@
Top 500 GPT apps. Creators of popular GPT apps are more likely to engage with trending information, which can lead them to adjust their GPT configurations. For example, they are likely to adopt and utilize superior, more functional system prompts from others. Analyzing these popular apps can better reveal the prevalence of plagiarism in GPT apps. Therefore, we investigate the similarity of system prompts among the top 500 GPT apps each week. We use the same approach to group similar GPT app responses, as the one used in analyzing the top 10,000 GPT apps. Additionally, we analyze responses from GPT apps where our GPT configuration extraction method failed, to infer whether these popular GPT apps use similar protective system prompts.
@@
Figure 19 shows the number of GPT apps that have at least one counterpart with similar system prompts (with a cosine similarity threshold of 0.95). Across various dates of GPT configuration extraction, an average of 19 GPT apps correlate with at least one similar GPT app. Of these, there is an average of 11.2 GPT apps that our approach failed to extract their system prompts. For example, on April 24, we found 14 GPT apps responding with the same string “You are a GPT. Your name is GPT” and 4 GPT apps replying “Sorry, bro! Not possible.”. This suggests a high likelihood of using similar protective system prompts. In terms of GPT apps that we successfully acquired their GPT configurations, a maximum of 11 GPT apps share nearly identical system prompts with different names in the March 13 data point. Further analysis reveals that these 11 GPT apps come from two GPT app creators; one creates 10 GPT apps and the other creates one. Additionally, two popular GPTs ranked 3rd and 8th in the same category had identical prompts, indicating a propensity for plagiarism among popular GPTs.
Takeaways: The ease of accessing GPT app configurations has led to a notable issue of plagiarism within the current GPT app store. It’s a significant issue in app ecosystems, impacting developer morale, reducing innovation, and potentially harming the user experience. It is crucial for GPT app stores to adopt some strategies to prevent this phenomenon. For example, implementing more rigorous review processes that thoroughly check new app submissions for similarities to existing apps can help prevent plagiarized apps from being listed. Besides, enabling users and developers to easily report suspected plagiarism is a good method to help app stores quickly identify and take action against such apps.
6. Recommendation
Based on our findings, we provide some recommendations for users, creators, and app stores to improve the LLM app ecosystem.
GPT App Users. Unlike traditional mobile apps that often provide detailed textual descriptions, UI designs, and screenshots to illustrate their functionality, many GPT apps lack comprehensive use-case descriptions that clarify what inputs produce specific outputs. Therefore, user ratings and feedback become crucial, especially in the initial stages of GPT stores and apps, as they guide potential users in making informed choices about which apps to use. We suggest users can actively provide user ratings.
GPT app creators. GPT app creators should be aware of the vulnerabilities inherent in GPT apps and make the protection of their configurations a priority. Given greater difficulties in acquiring external APIs, creators can transfer sensitive data to these APIs to prevent potential data leaks. Additionally, incorporating protective instructions within system prompts can further complicate unauthorized access to these prompts, enhancing security.
GPT app stores. App stores serve as crucial intermediaries between creators and users, playing a vital role in the development of the app market and ecosystem. From the perspective of dynamics and evolution, app stores could implement tangible incentives to foster engagement from both creators and users. Encouraging users to submit ratings and comments can facilitate the continuous development and refinement of apps. Similarly, providing financial incentives can spur creators’ creativity and innovation, motivating them to explore new ideas and enhance the functionality of their apps.
From the security side, app stores should implement robust measures to safeguard the integrity and privacy of the apps they host. For example, they could set restricted review processes to screen apps for vulnerabilities and potential security risks and conduct continuous monitoring to detect and respond to threats promptly. Additionally, implementing strict data protection protocols and security barriers can help prevent unauthorized access and data leaks, ensuring a safe and secure environment for both users and developers. These practices are crucial for maintaining user trust and fostering a secure and sustainable app ecosystem.
@@ 8. Conclusion
In this paper, we present the first large-scale measurement study of two GPT app stores, offering a preliminary analysis of the LLM app ecosystem. Utilizing our automated tools, we monitor these app stores over a period of five months and develop a comprehensive dataset, which contains three types of data: metadata, user feedback, and GPT configuration. Initially, we explore the landscape of GPT apps, examining their dynamics and evolution. We discover that while user enthusiasm for GPT apps continues to grow, creator interest tends to level off within three months of the apps’ introduction. Then, we investigate the vulnerabilities and potential security issues within app stores. We observe that nearly 90% of system prompts are readily accessible, indicating a widespread failure to secure app configurations, which leads to significant plagiarism and duplication among apps. Our findings emphasize the need for app stores, creators, and users to work together to improve the LLM app ecosystem.
@@ Appendix A Ethics
We confirm that this study does not raise any ethical issues from the following three aspects: First, this study is IRB-approved; Second, our study strictly complies with the policies in OpenAI Terms of Use (OpenAI, 2023). In particular, our data collection and analysis are conducted solely for research purposes; In addition, all our results and findings were obtained from the publicly accessible data. We also ensure that the collected dataset and analysis results will never be used by us for any illegal, harmful, or abusive activities at any time – past, present, or future.
Appendix B Evaluation of Configuration Extraction

### v3

@@ 6. Plagiarism in the Wild
The GPT app store is the primary interface for users to find apps. As depicted in Figure 12, OpenAI’s GPT store displays up to ten related search results with similar names or functions when users search for a specific keyword, raising concerns about plagiarism among GPT apps. In this section, we explore the impact of these vulnerabilities by addressing two questions: (1) Are similar app names and descriptions impacting user experience during searches? (2) What are the similarities in system prompts for apps with similar names and descriptions, and why might GPT app creators choose to maintain these similarities?
@@ 6.1. Similarities in Names and Descriptions
We analyze duplicate and similar GPT apps in the OpenAI GPT app store by creating a search bot that automatically interacts with the store, using the top 10,000 app names and descriptions. We then verify: (1) Whether the target GPT app and search results have identical names and descriptions; (2) Whether they share the same name and semantically similar descriptions. We measure semantic similarity by converting descriptions into embeddings using a BERT model (bert-base-multilingual-uncased) (Wolf et al., 2020) and calculating cosine similarity. If the value is $\geq$ 0.99, we treat them as similar apps.
@@
Initially, we analyze GPT apps with identical names and descriptions. Figure 13 (left) shows 135 apps have an identical counterpart, and 21 out of 10,000 have two identical matches. Only a small number of GPT apps have more than two identical duplicates. However, the requirement for identical names and descriptions is strict. By loosening it to include semantically similar descriptions, we identify more similar GPT apps. Figure 13 (right) shows that over 700 GPT apps have at least one similar counterpart. These findings suggest that users searching for a specific GPT app may encounter multiple apps with identical names and similar descriptions, which can confuse users due to the limited information available in OpenAI’s GPT app store search interface.
@@
Then, we examine the conversation counts of similar GPT apps, which indicate popularity in the GPT app store. Users often rely on these counts when choosing apps. As shown in Figure 14, we assess the absolute difference in conversation counts between the target GPT app and its counterparts and find that most similar apps have slight differences in conversation counts, with over 75% having less than a 100-count difference. This makes it harder for users to choose the best app. Adding user ratings could be a helpful solution.
@@ 6.2. Similarities in System Prompts
To analyze the similarities in system prompts among apps, we utilize OpenAI’s text-embedding-3-small text embedding model to generate system prompt embeddings. The thresholds for cosine similarity are set to 0.9 and 0.95.
@@
Figure 15. Similarity Analysis in System Prompts Among the Top 10,000 GPT Apps with Two Cosine Similarity Thresholds: 0.9 and 0.95.
Top 10,000 GPT apps. Figure 15 displays the number of GPT apps with 1–6 similar counterparts. We find that 91 apps have a counterpart with a similarity of 0.9, and 45 apps have potential duplicates at a 0.95 threshold, contributing to 1.2% and 0.6% of all GPTs, respectively. Our further validation from their metadata shows that none of these GPT apps share the same names, indicating that few creators replicate GPT apps to resemble existing ones. Then, we categorize these similarities into two types. (1) Apps created by the same author with minor modifications in names and prompts to fit slightly different use cases, such as translating to English versus Chinese. This likely stems from the lack of user-friendly options for switching functionality (e.g., setting the translation target to English or Chinese) in the UI of GPT apps, as user interactions with these apps are currently text-based. This significantly constrains user-friendliness for both app developers and users. (2) Apps created by different authors with distinct names. We infer that some creators mimic or plagiarize popular apps to attract more users and increase profit.
@@
Figure 16. Similarity Analysis in System Prompts Among the Top 500 GPT Apps. “Success” targets GPT apps successfully in GPT configuration extraction, whereas “Fail” targets the rest of the GPT apps.
@@
Top 500 GPT apps. Creators of popular apps often adapt to trends by adjusting GPT configurations, sometimes adopting more functional prompts from others. Analyzing popular apps can help reveal plagiarism among GPT apps. To investigate, we examine the similarity of system prompts among the top 500 GPT apps each week. We also analyze responses from GPT apps where our extraction method failed, to determine if they use similar protective system prompts.
@@
Figure 16 shows on average, 14 GPT apps have at least one counterpart with similar prompts (cosine similarity threshold of 0.95). Of these, an average of 7 apps per week failed to reveal their prompts. For example, on March 20, we found 17 apps responding identically with “You are a GPT. Your name is GPT.”, and 7 apps replying “Sorry, bro! Not possible.”, indicating the use of the same protective prompts. For apps where we successfully extract configurations, up to 20 GPT apps had nearly identical system prompts on July 31. Additionally, on March 20, two popular GPTs ranked 3rd and 8th in the same category also shared identical prompts, pointing to potential plagiarism among top apps.
$\bullet$ Findings: The ease of accessing GPT app configurations has led to significant plagiarism in the GPT app store, affecting developer morale, innovation, and user experience. To combat this, GPT app stores should implement stricter review processes to check for app similarities and allow easy reporting of suspected plagiarism to take prompt action.
@@ 7. Related Work
Measurement Study of GPT Apps. Several studies have discussed the current state of GPT stores and GPT apps; however, we are the first to provide a comprehensive longitudinal study lasting 10 months. Some research only analyzed GPT app metadata, such as app names, descriptions, and ratings (Zhao et al., 2024b, a). Zhao et al. (Zhao et al., 2024a) categorized potential risks faced by GPT apps but lacked real-world data for validation. Su et al. (Su et al., 2024) employed similar approaches to ours to extract system prompts from GPT apps, but they conducted small-scale experiments on fewer than 1,000 GPT apps—an order of magnitude smaller than our study. Another recent study (Hou et al., 2024) explored whether GPT apps generate malicious content or violate privacy policies. Yan et al. (Yan et al., 2024) examined the GPT app plugin subsystem. The aforementioned work is either covered by us or orthogonal to our research objectives. Additionally, our work extracts and analyzes GPT app configurations over time, providing a comprehensive quantification of the landscape and vulnerability of GPT apps in the wild.
@@ 8. Conclusion
In this paper, we present the first large-scale measurement study of three GPT app stores, analyzing the LLM app ecosystem over a ten-month period. Using automated tools, we gather the app’s information, including metadata, user feedback, and GPT configurations. We find that the ecosystem has stabilized with few new apps and creators, but app usage is imbalanced, with only a small fraction actively used. Additionally, nearly 90% of system prompts are easily accessible, leading to significant plagiarism and duplication. Thus, there is a need for app stores, creators, and users to collaborate in improving the LLM app ecosystem’s security and growth.
We suggest that users should actively provide ratings and feedback to improve app discoverability and guide potential users, especially since many apps lack detailed use-case descriptions. GPT app creators should prioritize security by transferring sensitive data to external APIs and incorporating protective instructions in system prompts to prevent unauthorized access. GPT app stores can implement financial incentives to boost engagement from both creators and users while also enforcing stronger security measures, including strict review processes and continuous monitoring, to detect vulnerabilities and protect user data.
@@ Appendix A Ethics
We confirm that this study does not raise any ethical issues. First, our study strictly complies with the policies in OpenAI Terms of Use (OpenAI, 2023). In particular, our data collection and analysis are conducted solely for research purposes. In addition, all our results and findings were obtained from publicly accessible data. We also ensure that the collected dataset and analysis results will never be used by us for any illegal, harmful, or abusive activities at any time – past, present, or future.
Appendix B GPT Store & App Statistics
B.1. Categories
Figure 18 shows the detailed names of 50 categories GPTStore.AI. OpenAI lists the top 16 apps in each category, each boasting over 100,000 conversation counts. Due to the app store’s practice of consistently featuring these popular apps, an increasing number of users are drawn to try them, thereby further elevating their conversation counts. However, this approach restricts visibility for other apps within the same category, which complicates user navigation and limits exposure to a broader array of app options.
B.2. Features and Functions
GPT apps are equipped with five major functions and various customized APIs to enhance their performance. Here is a description of these five key functions. File Attachment: Users can upload files to apps. File Browser: This function enables users to access the web during conversations. DALL·E: Image generation capabilities are enabled with DALL·E. Python: Apps have the capability to write and execute Python codes. Knowledge File: GPT apps contain knowledge files that store essential information,
Figure 18. The overview of 50 categories
Appendix C Evaluation of Configuration Extraction